# Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential

### v3

@@ Abstract
Introduction. How can classical behaviors of objects, which obey our intuition about how the macroscopic world behaves, be distinguished from non-classical behaviors such as what are described by quantum theory? The Leggett-Garg inequality (LGI) was proposed to answer this question LG ; Leggett ; FN . It serves as a test of a concept called macrorealism, which consists of two conditions: macrorealism per se (MRPS) and noninvasive measurability (NIM). MRPS assumes that a macroscopic system with two or more macroscopically distinct states available to it is always in one or the other of these states, and NIM assumes that it is possible to determine the state of the system with arbitrary small perturbations on its subsequent dynamics. A violation of the LGI implies a departure from macrorealism. The exploration of the LGI was inspired by the question of whether macroscopic coherence, as illustrated by Schrödinger’s cat gedanken experiment, can be realized in a laboratory cat .
@@
The violation of the LGI has been studied in several systems, experimentally and theoretically FN ; Emary ; Barbieri ; Avis ; Kofler ; Budroni_2013 ; Budroni ; Moreira ; Lambert ; Formaggio ; Martin ; Laura ; Matsumura ; Mawby . Recently, experimental LG tests have been achieved in a superconducting flux qubit with genuine macroscopicity, refuting its classical realistic description flux .
@@
The system of Bose-Einstein condensates (BECs) in a double-well potential is suitable for testing the LGI to clarify the boundary between classical and quantum regimes. The dynamics of a BEC in a double-well potential has been observed in several experiments Albiez ; Levy ; Gati ; Berrada ; Ryu ; Shin ; Saba ; Mukhopadhyay ; Vretenar . In particular, in the experiment of the Heidelberg group Albiez , Josephson oscillations and macroscopic quantum self-trapping were successfully observed, and the experimental result exhibits agreement with the prediction by the Gross-Pitaevskii (GP) equation Smerzi ; Gross ; Pitaevskii . It implies that a population imbalance and a phase difference between the two wells have definite values independent of measurements in the presence of sufficiently large number of particles in the double-well potential GP .
@@
In this Letter, we discuss the violation of the LGI for bosons in a double-well potential. We find the violation of the LGI in the BEC regime, in which the system can be considered as two weakly coupled BECs, despite the fact that their dynamics are well described by coupled semiclassical equations for definite phase and population differences, which are presumed to be unaffected by measurements. We reveal that the violation of the LGI is associated with the collapse of the state due to projective measurements in the intermediate time, which necessarily affects the subsequent dynamics. We find that the critical value of interaction strength for the onset of the violation of the LGI coincides with the one for self-trapping. Furthermore, we observe that the violation of the LGI becomes increasingly significant as the particle number increases.
In the single-particle regime, in which bosons do not interact with each other, the bosons tunneling backwards induced by the measurement are found to be crucial for the violation of the LGI.
@@
Model. The system of bosons in a double-well potential is well described by the Bose-Josephson junction (BJJ) model Smerzi ; Zapata ; Raghavan ; Ferrini ; Hipolito ; Xu ; Vardi ; Zibold ; Pezze ; Milburn ; Sinha ; Vardi2 . Using the Schwinger boson representation $\hat{S}_{x}=(\hat{a}_{\rm L}^{\dagger}\hat{a}_{\rm R}+\hat{a}_{\rm R}^{\dagger}\hat{a}_{\rm L})/2$ and $\hat{S}_{z}=(\hat{n}_{\rm L}-\hat{n}_{\rm R})/2$, where $\hat{a}_{\alpha}$ ($\hat{a}_{\alpha}^{\dagger}$) and $\hat{n}_{\alpha}=\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}$ are annihilation (creation) and number operators of a boson in the well $\alpha={\rm L},{\rm R}$, the Hamiltonian of the BJJ model can be written as Zibold ; Pezze ; Milburn
@@
Method.—The LGI is formulated for the correlation functions $C_{ij}=\langle Q_{i}Q_{j}\rangle$ of dichotomic variables $Q_{i}=\pm 1$, which is given as the result of the measurement at time $t_{i}$, where $i,j=1,2,3$. MRPS ensures the existence of the value of the observable $Q_{i}$ at $t_{i}$, regardless of whether or not the measurement is performed. Thus we can define the joint probabilities $P_{ij}(Q_{1},Q_{2},Q_{3})$, where measurements are performed at $t_{i}$ and $t_{j}$, whereas the system is unmeasured at time $t_{k}$ ($\neq t_{i},t_{j}$). Under only MRPS, the joint probability may depend on when measurements are made, since measurements at different times can affect the time-evolution differently. By adding the NIM condition, however, the subscripts $i$ and $j$ can be dropped since any measurement does not affect the system under NIM. Then the joint probability $P(Q_{1},Q_{2},Q_{3})$ yields all correlation functions and we obtain the LGI LG ; FN .
@@
BEC regime. First, we examine violation of the LGI in the BEC regime ($\Lambda\neq 0$ and $N\gg 1$), in which the system can be considered as two weakly coupled BECs. Such BECs have been studied by a semiclassical approach based on two coupled GP equations, which predict Josephson oscillation or self-trapping depending on the value of $\Lambda$ and population imbalance at initial time Albiez ; Levy ; Smerzi ; Milburn ; Julia ; Ananikian .
@@
Population dynamics in the BEC regime. Figures 11 present the time-evolutions of the occupation probability of the left well $P(n_{\rm L})$ in the case of $N=2000$ with the initial state $|N\rangle_{\rm L}|0\rangle_{\rm R}$. The peak of the occupation probability oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ due to the single-particle tunneling effect when $\Lambda=0$ [Fig. 1] A similar coherent oscillation occurs in the presence of an on-site interaction, but it is rather induced by the Josephson effect when $0<\Lambda<\Lambda_{\rm st}$ [Figs. 11]. For a strong interaction strength, the Josephson oscillation ceases and self-trapping occurs when $\Lambda>\Lambda_{\rm st}$, where the peak of the occupation probability oscillates keeping the population imbalance [Fig. 1]. The critical value of interaction strength for the onset of self-trapping $\Lambda_{\rm st}\simeq 2$ agrees well with the value derived in the semiclassical analysis based on the GP equation $\Lambda_{\rm st}^{\rm GP}=2$ Smerzi .
@@
If the dynamics obeys coupled GP equations, oscillation of the population imbalance continues without damping Smerzi ; Raghavan ; Ananikian . The damping of the oscillation that occurs in the presence of an on-site interaction in Figs. 11 can be understood as being due to the intrinsic quantum fluctuation in the initial state Milburn . That is, since the initial population imbalance is fixed, the initial phase difference between the BECs should be fluctuating due to the uncertainty relation between the population imbalance and phase difference Milburn . As $n_{\rm L}$ and $n_{\rm R}$ becomes larger, the quantum fluctuation ($\simeq 1/\sqrt{n_{{\rm L},{\rm R}}}$ Gajda ) becomes more negligible Pezze ; Smerzi . This is the reason why, with the increase of $N$, the coherent Josephson oscillation persists for longer (see Fig. S3 in the SM).
@@
BECs in a double-well potential are considered to possess macroscopic observables for sufficiently large $N$. In fact, the damping is suppressed for larger $N$, and the population dynamics is consistent with the semiclassical description (see SM). The experimental data in Ref. Albiez have also been shown to be consistent with the prediction of the semiclassical analysis. Therefore, both the theoretical and experimental studies support the existence of macroscopic wave functions of two BECs with a definite phase difference and population imbalance for sufficiently large $N$. Upon such considerations, it seems reasonable to expect nonviolation of the LGI.
@@
Violation of the LGI in the BEC regime. Figure 1 shows the result of the LG test in the BEC regime for $N=2000$: $K$ is plotted as a function of $\tau$ and $\Lambda$. $K$ exhibits periodic oscillations as a function of $\tau$ reflecting the oscillation of the occupation probabilities in Figs. 11. Contrary to the expectation, the LGI is violated near the peaks of the oscillations for the interaction $-\Lambda_{\rm c}<\Lambda<\Lambda_{\rm c}$. For $|\Lambda|\geq\Lambda_{\rm c}$ the LGI is nonviolated. Remarkably, the critical value $\Lambda_{\rm c}$ that separates violation and nonviolation of the LGI coincides with $\Lambda_{\rm st}$ as illustrated in Fig. 1. Thus, the LGI is violated due to Josephson oscillations when $\Lambda\neq 0$, whereas at $\Lambda=0$, the violation is due to single-particle oscillations. The violation of the LGI for $\Lambda\neq 0$ implies that macrorealism, which consists of MRPS and NIM, is excluded even for such a large $N$, and thus contradicts what is expected by the GP description; the LGI is never broken for the GP description GP . In Figs. 1 and 1, the LGI is also violated at smaller values of $N$ for $|\Lambda|$ below the critical values near 2. These critical values coincide with the thresholds for the onset of self-trapping for respective $N$.
@@
Figure 1 shows the maximal violations of the LGI for particle numbers corresponding to Figs. 11. It indicates that the LGI is no longer violated, regardless of $\tau$, once $\Lambda$ exceeds the critical value $\Lambda_{\rm c}(N)$ for each $N$. Notably, this critical value approaches the threshold for the onset of self-trapping, but remains below, $\Lambda_{\rm st}^{\rm GP}=2$ as $N$ increases. Remarkably, the violation of the LGI becomes more significant as $N$ increases for fixed $\Lambda$, as shown in Figs. 1 and 1 (also see SM).
Zárate et al. investigated the violation of the LGI for two weakly coupled BECs in a double-well potential and found a maximal violation of the LGI 1.5 Laura . This is attributed to their presumption of the two-state approximation that results in two-outcome measurements Budroni_2013 (also see SM). In contrast, Fig. 1 shows an LG violation exceeding 2, consistent with the result in Ref. Budroni , due to the use of multiple projectors |$\ket{n_{\rm L}}\ket{n_{\rm R}}$ and coarse graining of the results by the sign ($\hat{S}_{z}$).
[…]
Figures 1 and 1 show that the period of oscillation of $K$ increases and its amplitude decreases as $\Lambda$ approaches $\Lambda_{\rm c}(N)$ from below. It is due to the softening of the Josephson oscillation associated with the transition from the Josephson regime to the self-trapping regime Pezze ; Farolfi .
@@
We numerically test the LGI for nearly the same parameters for which Josephson oscillations were observed in Ref. Albiez . Here, we set $\Lambda=15$ and $N=1000$. The LGI is shown to be violated for realistic parameters in Fig. 3, in which the evolution of $K$ is plotted as a function of $\tau$ for several initial population imbalances $z(0)=(n_{\rm L}(0)-n_{\rm R}(0))/N$. $z(0)$ needs to be smaller than the critical value for self-trapping $z_{\rm st}(0)\simeq 0.5$ Raghavan . $z(0)=0.28$ is expected to be more desirable for the demonstration of violation of the LGI. Figure 3 shows the time evolution of $P(n_{\rm L})$ for $z(0)=0.28$. It exhibits the tunneling of atoms between the two wells, which indeed leads to the violation of the LGI.
@@
The violation of the LGI in Fig. 4 closely resembles the one presented in Fig. 2(a) in Ref. Lambert . This is due to the special coarse graining of the results as the sign ($S_{z}$). Our work represents a concrete realization of the proposal of Lambert et al. for the coarse-grained implementation of LGI violations.
@@
Experimental realization. The experimental setup in Ref. Albiez provides a concrete realization of the theoretical predictions in our paper. Phase contrast imaging (PCI) is a promising candidate for implementing our measurement scheme as a nondestructive measurement of a BEC in a double-well potential Andrews ; Bradley ; Higbie ; Ilo-Okeke . PCI can perform measurements in the number basis |$\ket{n_{\rm L}}\ket{n_{\rm R}}$, enabling the measurement of the population in each well nondestructively without atom loss. This approach allows for the acquisition of each correlation function.
Conclusion and discussion. In this Letter, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak-coupling regime, while it remains unviolated in the strong-coupling regime for $\Lambda$ beyond the critical $\Lambda_{\rm c}$, which depends on the particle number $N$. $\Lambda_{\rm c}$ approaches, but remains below, the critical threshold for the onset of self-trapping, $\Lambda_{\rm st}^{\rm GP}=2$, as derived from semiclassical analysis. It indicates that the nonviolation of the LGI is attributed to the self-trapping phenomena. Remarkably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. We also find that, in this system, an on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. Furthermore, we have revealed that the LGI is violated for nearly the same parameters for which the Josephson oscillations are observed in Ref. Albiez , thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.

### v5

@@ Abstract
Introduction. How can classical behaviors of objects, which obey our intuition about how the macroscopic world behaves, be distinguished from non-classical behaviors such as what are described by quantum theory? The Leggett-Garg inequality (LGI) was proposed to answer this question [1, 2, 3]. It serves as a test of a concept called macrorealism, which consists of two conditions: macrorealism per se (MRPS) and noninvasive measurability (NIM). MRPS assumes that a macroscopic system with two or more macroscopically distinct states available to it is always in one or the other of these states, and NIM assumes that it is possible to determine the state of the system with arbitrary small perturbations on its subsequent dynamics. A violation of the LGI implies a departure from macrorealism. The exploration of the LGI was inspired by the question of whether macroscopic coherence, as illustrated by Schrödinger’s cat gedanken experiment, can be realized in a laboratory [4].
@@
The violation of the LGI has been studied in several systems, experimentally and theoretically [3, 7, 5, 6, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Recently, experimental LG tests have been achieved in a superconducting flux qubit with genuine macroscopicity, refuting its classical realistic description [18].
@@
The system of Bose-Einstein condensates (BECs) in a double-well potential is suitable for testing the LGI to clarify the boundary between classical and quantum regimes. The dynamics of a BEC in a double-well potential has been observed in several experiments [21, 23, 22, 24, 25, 19, 20, 27, 26]. In particular, in the experiment of the Heidelberg group [21], Josephson oscillations and macroscopic quantum self-trapping were successfully observed, and the experimental result exhibits agreement with the prediction by the Gross-Pitaevskii (GP) equation [30, 29, 28]. It implies that a population imbalance and a phase difference between the two wells have definite values independent of measurements in the presence of sufficiently large number of particles in the double-well potential [31].
@@
In this Letter, we discuss the violation of the LGI for bosons in a double-well potential. We find the violation of the LGI in the BEC regime, in which the system can be considered as two weakly coupled BECs, despite the fact that their dynamics are well described by coupled semiclassical equations for definite phase and population differences, which are presumed to be unaffected by measurements. We reveal that the violation of the LGI is associated with the collapse of the state due to projective measurements in the intermediate time, which necessarily affects the subsequent dynamics. We find that the critical value of interaction strength for the onset of the violation of the LGI coincides with the one for self-trapping. Furthermore, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. In the single-particle regime, in which bosons do not interact with each other, the bosons tunneling backwards induced by the measurement are found to be crucial for the violation of the LGI.
@@
Model. The system of bosons in a double-well potential is well described by the Bose-Josephson junction (BJJ) model [30, 35, 33, 37, 36, 39, 42, 38, 40, 32, 41, 34]. Using the Schwinger boson representation $\hat{S}_{x}=(\hat{a}_{\rm L}^{\dagger}\hat{a}_{\rm R}+\hat{a}_{\rm R}^{\dagger}\hat{a}_{\rm L})/2$ and $\hat{S}_{z}=(\hat{n}_{\rm L}-\hat{n}_{\rm R})/2$, where $\hat{a}_{\alpha}$ ($\hat{a}_{\alpha}^{\dagger}$) and $\hat{n}_{\alpha}=\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}$ are annihilation (creation) and number operators of a boson in the well $\alpha={\rm L},{\rm R}$, the Hamiltonian of the BJJ model can be written as [38, 40, 32]
@@
Method.—The LGI is formulated for the correlation functions $C_{ij}=\langle Q_{i}Q_{j}\rangle$ of dichotomic variables $Q_{i}=\pm 1$, which is given as the result of the measurement at time $t_{i}$, where $i,j=1,2,3$. MRPS ensures the existence of the value of the observable $Q_{i}$ at $t_{i}$, regardless of whether or not the measurement is performed. Thus we can define the joint probabilities $P_{ij}(Q_{1},Q_{2},Q_{3})$, where measurements are performed at $t_{i}$ and $t_{j}$, whereas the system is unmeasured at time $t_{k}$ ($\neq t_{i},t_{j}$). Under only MRPS, the joint probability may depend on when measurements are made, since measurements at different times can affect the time-evolution differently. By adding the NIM condition, however, the subscripts $i$ and $j$ can be dropped since any measurement does not affect the system under NIM. Then the joint probability $P(Q_{1},Q_{2},Q_{3})$ yields all correlation functions and we obtain the LGI [1, 3].
@@
BEC regime. First, we examine violation of the LGI in the BEC regime ($\Lambda\neq 0$ and $N\gg 1$), in which the system can be considered as two weakly coupled BECs. Such BECs have been studied by a semiclassical approach based on two coupled GP equations, which predict Josephson oscillation or self-trapping depending on the value of $\Lambda$ and population imbalance at initial time [21, 23, 30, 32, 44, 43].
@@
Population dynamics in the BEC regime. Figures 11 present the time-evolutions of the occupation probability of the left well $P(n_{\rm L})$ in the case of $N=2000$ with the initial state $|N\rangle_{\rm L}|0\rangle_{\rm R}$. The peak of the occupation probability oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ due to the single-particle tunneling effect when $\Lambda=0$ [Fig. 1] A similar coherent oscillation occurs in the presence of an on-site interaction, but it is rather induced by the Josephson effect when $0<\Lambda<\Lambda_{\rm st}$ [Figs. 11]. For a strong interaction strength, the Josephson oscillation ceases and self-trapping occurs when $\Lambda>\Lambda_{\rm st}$, where the peak of the occupation probability oscillates keeping the population imbalance [Fig. 1]. The critical value of interaction strength for the onset of self-trapping $\Lambda_{\rm st}\simeq 2$ agrees well with the value derived in the semiclassical analysis based on the GP equation $\Lambda_{\rm st}^{\rm GP}=2$ [30].
@@
If the dynamics obeys coupled GP equations, oscillation of the population imbalance continues without damping [30, 33, 43]. The damping of the oscillation that occurs in the presence of an on-site interaction in Figs. 11 can be understood as being due to the intrinsic quantum fluctuation in the initial state [32]. That is, since the initial population imbalance is fixed, the initial phase difference between the BECs should be fluctuating due to the uncertainty relation between the population imbalance and phase difference [32]. As $n_{\rm L}$ and $n_{\rm R}$ becomes larger, the quantum fluctuation ($\simeq 1/\sqrt{n_{{\rm L},{\rm R}}}$ [45]) becomes more negligible [40, 30]. This is the reason why, with the increase of $N$, the coherent Josephson oscillation persists for longer (see Fig. S3 in the SM).
@@
BECs in a double-well potential are considered to possess macroscopic observables for sufficiently large $N$. In fact, the damping is suppressed for larger $N$, and the population dynamics is consistent with the semiclassical description (see SM). The experimental data in Ref. [21] have also been shown to be consistent with the prediction of the semiclassical analysis. Therefore, both the theoretical and experimental studies support the existence of macroscopic wave functions of two BECs with a definite phase difference and population imbalance for sufficiently large $N$. Upon such considerations, it seems reasonable to expect nonviolation of the LGI.
@@
Violation of the LGI in the BEC regime. Figure 1 shows the result of the LG test in the BEC regime for $N=2000$: $K$ is plotted as a function of $\tau$ and $\Lambda$. $K$ exhibits periodic oscillations as a function of $\tau$ reflecting the oscillation of the occupation probabilities in Figs. 11. Contrary to the expectation, the LGI is violated near the peaks of the oscillations for the interaction $-\Lambda_{\rm c}<\Lambda<\Lambda_{\rm c}$. For $|\Lambda|\geq\Lambda_{\rm c}$ the LGI is nonviolated. Remarkably, the critical value $\Lambda_{\rm c}$ that separates violation and nonviolation of the LGI coincides with $\Lambda_{\rm st}$ as illustrated in Fig. 1. Thus, the LGI is violated due to Josephson oscillations when $\Lambda\neq 0$, whereas at $\Lambda=0$, the violation is due to single-particle oscillations. The violation of the LGI for $\Lambda\neq 0$ implies that macrorealism, which consists of MRPS and NIM, is excluded even for such a large $N$, and thus contradicts what is expected by the GP description; the LGI is never broken for the GP description [31]. In Figs. 1 and 1, the LGI is also violated at smaller values of $N$ for $|\Lambda|$ below the critical values near 2. These critical values coincide with the thresholds for the onset of self-trapping for respective $N$.
@@
Figure 1 shows the maximal violations of the LGI for particle numbers corresponding to Figs. 11. It indicates that the LGI is no longer violated, regardless of $\tau$, once $\Lambda$ exceeds the critical value $\Lambda_{\rm c}(N)$ for each $N$. Notably, this critical value approaches, but remains below, the threshold for the onset of self-trapping $\Lambda_{\rm st}^{\rm GP}=2$ as $N$ increases. Remarkably, the violation of the LGI becomes more significant as $N$ increases for fixed $\Lambda$, as shown in Figs. 1 and 1 (also see SM).
Zárate et al. investigated the violation of the LGI for two weakly coupled BECs in a double-well potential and found a maximal violation of the LGI 1.5 [15]. This is attributed to their presumption of the two-state approximation that results in two-outcome measurements [9] (also see SM). In contrast, Fig. 1 shows an LG violation exceeding 2, consistent with the result in Ref. [10], due to the use of multiple projectors |$\ket{n_{\rm L}}\ket{n_{\rm R}}$ and coarse graining of the results by the sign ($\hat{S}_{z}$).
Figures 1 and 1 show that the period of oscillation of $K$ increases and its amplitude decreases as $\Lambda$ approaches $\Lambda_{\rm c}(N)$ from below. It is due to the softening of the Josephson oscillation associated with the transition from the Josephson regime to the self-trapping regime [40, 46].
@@
We numerically test the LGI for nearly the same parameters for which Josephson oscillations were observed in Ref. [21]. Here, we set $\Lambda=15$ and $N=1000$. The LGI is shown to be violated for realistic parameters in Fig. 3, in which the evolution of $K$ is plotted as a function of $\tau$ for several initial population imbalances $z(0)=(n_{\rm L}(0)-n_{\rm R}(0))/N$. $z(0)$ needs to be smaller than the critical value for self-trapping $z_{\rm st}(0)\simeq 0.5$ [33]. $z(0)=0.28$ is expected to be more desirable for the demonstration of violation of the LGI. Figure 3 shows the time evolution of $P(n_{\rm L})$ for $z(0)=0.28$. It exhibits the tunneling of atoms between the two wells, which indeed leads to the violation of the LGI.
@@
The violation of the LGI in Fig. 4 closely resembles the one presented in Fig. 2(a) in Ref. [12]. This is due to the special coarse graining of the results as the sign ($S_{z}$). Our work represents a concrete realization of the proposal of Lambert et al. for the coarse-grained implementation of LGI violations.
@@
Experimental realization. The experimental setup in Ref. [21] provides a concrete realization of the theoretical predictions in our paper. Phase contrast imaging (PCI) is a promising candidate for implementing our measurement scheme as a nondestructive measurement of a BEC in a double-well potential [47, 48, 49, 50]. PCI can perform measurements in the number basis |$\ket{n_{\rm L}}\ket{n_{\rm R}}$, enabling the measurement of the population in each well nondestructively without atom loss. This approach allows for the acquisition of each correlation function.
Conclusion and discussion. In this Letter, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak-coupling regime, while it remains unviolated in the strong-coupling regime for $\Lambda$ beyond the critical $\Lambda_{\rm c}$, which depends on the particle number $N$. $\Lambda_{\rm c}$ approaches, but remains below, the critical threshold for the onset of self-trapping, $\Lambda_{\rm st}^{\rm GP}=2$, as derived from semiclassical analysis. It indicates that the nonviolation of the LGI is attributed to the self-trapping phenomena. Remarkably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. We also find that, in this system, an on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. Furthermore, we have revealed that the LGI is violated for nearly the same parameters for which the Josephson oscillations are observed in Ref. [21], thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.